# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v2

@@ Appendix B Proof of (5)
$+\mathbb{E}_{p_{D}}\left[\frac{g^{2}(t)}{2}\left(\left\|\nabla\log p_{D}(z_{t})\right\|^{2}+\left\|s\left(z_{t},x,t|\theta\right)\right\|^{2}\right.\right.$
@@
$\quad\quad\quad\quad\ \quad\quad\left.\left.-2s\left(z_{t},x,t|\theta\right)^{\top}\nabla\log p_{D}(z_{t}|z_{0})\right)\right].$
Appendix C Implementation of decoder
The detailed implementation of the GHD based decoder is summarized in Algorithm 1. Within the algorithm, we initially stochastically generate a sample $y$ using the random variable $\zeta$ (refer to Line 1 of the algorithm). Subsequently, we iteratively update the sample using random velocities $v_{1},\ldots,v_{K}$ and GHD (see Line 5). Finally, we utilize discretized Brownian dynamics with a trainable step size to design the decoder density of $x$ (refer to Line 9). Here, $\epsilon(l;\phi)$, $Q_{v}$, $Q_{y}$, $T_{v}$ and $T_{y}$ are all neural networks composed of three-layer MLPs. $\epsilon_{0}$ serves as a hyperparameter, and our numerical experiments suggest that setting $\epsilon_{0}$ to a small positive value can improve the stability of the algorithm.
@@ Appendix C Implementation of decoder
1:$z_{0}=(\zeta,v_{1},\ldots,v_{K})$ drawn from the standard Gaussian distribution and decoder parameters $\phi$
@@
6:         Updata $(y,v_{i})$ by
$v_{k}$ ::=:= $v_{k}-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right)$
$y$ ::=:= $y+\epsilon(l;\phi)\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},l;\phi)}+T_{y}(v_{k},l;\phi)\right)$
$v_{k}$ ::=:= $v_{k}-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right)$
@@
10:Let $\mu=y-e^{\epsilon_{0}\eta(y;\phi^{\prime})}\nabla U(y)$ and $\Sigma=2e^{\epsilon_{0}\eta(y;\phi^{\prime})}I$.
11:return $x\sim\mathcal{N}(\mu,\Sigma)$.
Appendix D Importance weighting for accurate time integral
It can be seen from (7) that an ideal choice for the time proposal $p(t)$ in Monte Carlo estimation of the integral on the right-hand side is given by
$p(t)\propto g^{2}(t)\mathbb{E}_{p(\epsilon)p_{D}(x,z_{t})}\left[\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)\right].$
In our experiments, we approximate the ideal proposal using the strategy introduced in [49]. The implementation details are as follows: We maintain a history buffer containing values of $\mathcal{L}_{t}$ from the most recent $30$ minibatches. In each iteration, the buffer provides an estimate of the expected value of $\mathcal{L}_{t}$ as a function of $t$ using a histogram with $100$ bins over the time interval $[0,T]$. Subsequently, we obtain an estimate of the ideal $p(t)$ for each iteration. Initially, during the early stages of training, $p(t)$ is set as a uniform distribution until $B_{\mathrm{buffer}}$ minibatches are accumulated.
Appendix E Experimental details
EDG In our experiments, we leverage the subVP SDE proposed in [37] to model the diffusion process of $z_{t}$ in (2), defined as
@@ Appendix E Experimental details
where $\zeta=(\zeta_{0},\zeta_{1})$, and $\mu_{0}(\zeta_{0};\phi),\Sigma_{0}(\zeta_{0};\phi)$ are modeled by MLPs. When incorporating the GHD in the decoder, we set $K=5$ and $J=5$.
Baseline For V-HMC, we facilitate the generation of samples after a warm-up phase comprising 1,000 steps, where each step the Metropolis-Hastings acceptance probability is computed to ensure the convergence. For L2HMC, we use a three-layer fully connected network, setting the leapfrog step length to 10. For different tasks, we adaptively select varying step sizes. For PIS, we use the publicly available code on our tasks directly. Regarding BG, the RealNVP architecture consists of 3 affine blocks, where the scaling and translation functions are modeled by a three-layer fully connected network, each with 256 units and ReLU activation functions.
@@
      fc $10\times 32$, ReLU, 2×2$\times$2 ×(fc $32\times 32$, ReLU), fc $32\times 2$.       fc $25\times 16$, ReLU,
@@
Final network $N(\mu,\Sigma)$:
@@
Given the reference samples $X=\{x_{i}\}_{i=1}^{m}$ and the generated samples $Y=\{y_{i}\}_{i=1}^{m}$, the Maximum Mean Discrepancy (MMD) [56] measures the difference between the distributions of $X$ and $Y$ as follows:
$\mathrm{MMD}^{2}(X,Y)=\frac{1}{m(m-1)}\sum_{i}\sum_{j\neq i}k(x_{i},x_{j})-\frac{2}{m^{2}}\sum_{i}\sum_{j}k(x_{i},y_{j})+\frac{1}{m(m-1)}\sum_{i}\sum_{j\neq i}k(y_{i},y_{j})$
where $k(\cdot,\cdot)$ denotes the kernel function to compute the inner product. RBF kernel is used in Table 1, and the bandwidth is set as the median distance between corresponding samples. We execute 20 cycles for each model to procure 5,000 desired samples, following which we evaluate the mean value of MMD in relation to 5,000 reference data. The standard deviations are all small, hence we do not report them in the table.
@@
      fc $d\times 256$, ReLU, 2×2$\times$2 ×(fc $256\times 256$, ReLU), fc $256\times 256$.       fc $(12d+1)\times 256$, ReLU,
@@
Final network $N(\mu,\Sigma)$:
@@
Ising model The model architecture is shown below, which is similar to that in NeuralRG [17] without stacking bijectors to form a reversible transformation. It retains the multiscale entanglement renormalization ansatz structure, while we only use one block to update the whole dimensions of the variable directly.
@@
      hierarchy network as shown in [17]       hierarchy network.
@@
Final network $N(\mu,\Sigma)$:
@@
Regarding the $p_{E}(z_{0}|x,\theta^{*})$, we employ the RK45 ODE solver implemented in scipy.integrate.solve_ivp, where the parameters set identical to those in [37]. The normalizing constant $\log Z$ is approximated according to Sec. 3.3.
Appendix F Ablation study
To elucidate the function of each component in EDG, we compare the following models on the sampling task of 2D energy functions: a VAE with the Gaussian decoder and encoder, where the means and diagonal covariance matrices are both parameterized by MLPs; a VAE with a GHD-based decoder and MLP-based encoder; EDG without GHD, where the decoder is modeled by MLPs; and the full EDG model. For VAE w/o GHD, the network is composed of 3-layer MLPs, each with 32 units and ReLU activation function. Starting from $z\sim N(0,I_{d}),d=10$, the decoder generates samples $x\sim p_{D}(x;\mu(z;\phi),\Sigma(z;\phi))$. The encoder, as an independent network, follows $p_{E}(z|x)=N(z;\mu(x;\theta),\Sigma(x;\theta))$. The objective function is the KL divergence of the joint distribution of $z$ and $x$ as

### v3

@@ Appendix C Proof of Theorem 1
$\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\left[\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\nabla_{z_{t}}\cdot s\left(z_{t},x,t;\theta\right)\right]\mathrm{d}t$
$=$ $\mathbb{E}_{t\sim\mathcal{U}[0,T],\epsilon\sim p(\epsilon),\left(x,z_{t}\right)\sim p_{D}\left(x,z_{t}\right)}\left[\frac{Tg(t)^{2}}{2}\left(\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\frac{\partial\left[\epsilon^{\top}s\left(z_{t},x,t;\theta\right)\right]}{\partial z_{t}}\epsilon\right)\right],$
where $\mathcal{U}[0,T]$ denotes a uniform distribution over $[0,T]$, and $p(\epsilon)\sim\mathrm{Rademacher}^{d}$ denotes a random vector in $\mathbb{R}^{d}$ with i.i.d. Rademacher entries, satisfying $\mathbb{E}[\epsilon]=0$ and $\mathrm{Cov}(\epsilon)=I$.
Appendix E Calculation of the probability flow ODE
Considering the SDE in Eq. (4), there is a corresponding deterministic ODE whose trajectories share the same marginal probability density
$\mathrm{d}z_{t}=\underbrace{\left\{f(z_{t},t)-\frac{1}{2}g(t)^{2}s_{\theta}(z_{t},x,t)\right\}}_{=:F_{\theta}(z_{t},x,t)}\mathrm{d}t.$
With the change of variables formula [ode], we can compute the log-likelihood of $p_{E}(z_{0}|x)$ using
$\log p_{E}(z_{0}|x)=\log p_{E}(z_{T})+\int_{0}^{T}\nabla\cdot F_{\theta}(z_{t},x,t)\mathrm{d}t.$
Due to the expensive computational cost of $\nabla\cdot F_{\theta}(z_{t},x,t)$, the Hutchinson trace estimator technique in D is used again to calculate
@@ Appendix E Calculation of the probability flow ODE
$\nabla\cdot F_{\theta}(z_{t},x,t)=\mathbb{E}_{p(\epsilon)}\left[\frac{\partial[\epsilon^{\top}F_{\theta}(z_{t},x,t)]}{\partial z_{t}}\epsilon\right],$
where $p(\epsilon)\sim\mathrm{Rademacher}^{d}$.
In our experiments, we employ the RK45 ODE solver implemented in scipy.integrate.solve_ivp, where the parameters set identical to those in [key-2] with atol=1e-5 and rtol=1e-5.
Appendix F Analysis of Eq. (14)
Based on Eq. (1) and $w(x,z_{0})=\frac{\exp(-U(x))p_{E}(z_{0}|x)}{p_{D}(z_{0})p_{D}(x|z_{0})}$, we have
$\log Z$ $\geq$ $D_{\mathrm{KL}}(p_{D}(x)\|\pi(x))-\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}\left[\log\frac{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}{\exp(-U(x))p_{E}(z_{0}|x;\theta)}\right]$
$=$ $D_{\mathrm{KL}}(p_{D}(x)\|\pi(x))+\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}\left[\log w(x,z_{0})\right].$
Therefore, any Monte Carlo estimator of $\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}\left[\log w(x,z_{0})\right]$ yields a lower bound on $\log Z$.
Moreover, according to A, as the model parameters are optimized during training, the distribution $p_{E}(z_{0}|x;\theta)$ is expected to approximate the true conditional distribution $p_{D}(z_{0}|x)$ more closely. This improved approximation deduces
$\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}\left[\log w(x,z_{0})\right]\rightarrow\log Z-D_{\mathrm{KL}}(p_{D}(x)\|\pi(x)).$
Consequently, a larger $\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}\left[\log w(x,z_{0})\right]$ suggests that the model is learning effectively, as $D_{\mathrm{KL}}(p_{D}(x)\|\pi(x))$ diminishes and the estimated $\log Z$ becomes more accurate.
@@
[…]
@@
6:   Update $(y,v_{i})$ by
$v_{k}$ $:=$ $v_{k}-\frac{\epsilon_{0}e^{\epsilon_{0}\epsilon(l;\phi)}}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right)$
$y$ $:=$ $y+\epsilon_{0}e^{\epsilon_{0}\epsilon(l;\phi)}\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},l;\phi)}+T_{y}(v_{k},l;\phi)\right)$
$v_{k}$ $:=$ $v_{k}-\frac{\epsilon_{0}e^{\epsilon_{0}\epsilon(l;\phi)}}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right)$
@@
10:Let $\mu=y-\epsilon_{0}e^{\epsilon_{0}\eta(y;\phi)}\nabla U(y)$ and $\Sigma=2\epsilon_{0}e^{\epsilon_{0}\eta(y;\phi)}I$.
11:return $x\sim\mathcal{N}(\mu,\Sigma)$, $p_{D}(x|z_{0};\phi)$.
Appendix G Implementation of decoder
The detailed implementation of the GHD based decoder is summarized in Algorithm 3. Within the algorithm, we initially stochastically generate a sample $y$ using the random variable $\zeta$ (refer to Line 1 of the algorithm). Subsequently, we iteratively update the sample using random velocities $v_{1},\ldots,v_{K}$ and GHD (see Line 5). Finally, we utilize discretized Brownian dynamics with a trainable step size to design the decoder density of $x$ (refer to Line 9). Here, $\epsilon$, $Q_{v}$, $Q_{y}$, $T_{v}$, $T_{y}$ and $\eta$ are all neural networks composed of three-layer MLPs. $\epsilon_{0}$ serves as an initial hyperparameter of trainable step size $\epsilon$ and $\eta$, and our numerical experiments suggest that setting $\epsilon_{0}$ to a small positive value can improve the stability of the algorithm.
Appendix H Experimental details
EDG In our experiments, we leverage the subVP SDE proposed in [key-2] to model the diffusion process of $z_{t}$ in Eq. (2), defined as
@@ Appendix H Experimental details
where $\zeta=(\zeta_{0},\zeta_{1})$, and $\mu_{0}(\zeta_{0};\phi),\Sigma_{0}(\zeta_{0};\phi)$ are modeled by MLPs. In the GHD decoder design, we fix $J=5$ while setting $M=2$ for 2D energy function tasks, $M=10$ for Bayesian Logistic Regression tasks and Ising models, and $M=100$ for Lennard-Jones.
Baseline For V-HMC, we facilitate the generation of samples after a warm-up phase comprising 1,000 steps, where each step the Metropolis-Hastings acceptance probability is computed to ensure the convergence. For L2HMC, we use a three-layer fully connected network, setting the leapfrog step length to 10. For different tasks, we adaptively select varying step sizes. For PIS, we use the publicly available code on our tasks directly. Regarding BG, the RealNVP architecture consists of 3 affine blocks, where the scaling and translation functions are modeled by a three-layer fully connected network, each with 256 units and ReLU activation functions. We would like to highlight that unlike data-driven energy-based models [chao2023training, nijkamp2020mcmc], our baseline BG generates samples solely from a prescribed energy function without any access to real data. The loss function is the KL divergence between the generated samples by BG and the target.
@@
   fc $10\times 32$, ReLU, $2\times$(fc $32\times 32$, ReLU), fc $32\times 2$.    fc $13\times 16$, ReLU,
@@
Final network ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}(\mu,\Sigma)$:
@@
Figure 5: Reweighting effect in Mog2(i) task. Left: The energy histograms of $1000$ generated samples under EDG and EDG (reweighted). Right: The mean value of the reweighted MMD with respect to the sample size. We report the mean value $\pm$ standard deviation over 10 runs.
@@
Given the reference samples $X=\{x_{i}\}_{i=1}^{m}$ with weights $w^{X}=\{w_{i}^{X}\}_{i=1}^{m}=\frac{1}{m}$ and the generated samples $Y=\{y_{i}\}_{i=1}^{m}$ with weights $w^{Y}=\{w_{i}^{Y}\}_{i=1}^{m}$, the Maximum Mean Discrepancy (MMD) [mmd] measures the difference between the distributions of $X$ and $Y$ as follows:
$\mathrm{MMD}^{2}(X,Y)=\sum_{i}\sum_{j}w_{i}^{X}w_{j}^{X}k(x_{i},x_{j})-2\sum_{i}\sum_{j}w_{i}^{X}w_{j}^{Y}k(x_{i},y_{j})+\sum_{i}\sum_{j}w_{i}^{Y}w_{j}^{Y}k(y_{i},y_{j}),$
where $k(\cdot,\cdot)$ denotes the kernel function to compute the inner product. RBF kernel is used in Table 1, and the bandwidth is set as the median distance between corresponding samples. We conduct 10 independent runs to generate 5,000 samples with uniform density, and compute the mean MMD against 5,000 reference samples. The standard deviations are all small, hence we do not report them in the table.
To better illustrate the effectiveness of the reweighting procedure in Sec. 3.3, we use Mog2(i)—a case with room for further improvement—as an example. Firstly, we display the energy histograms of the generated samples under EDG and EDG (reweighted) in the left panel of Fig. 5. Then, we compute the MMD using generated samples by EDG with normalized weights, which are calculated according to Sec. 3.3. The mean value across 10 independent runs are reported in Table 5, each based on 5,000 samples. The standard deviations are omitted as they are negligibly small. The convergence of the weighted MMD with respect to the sample size is also presented in the right panel of Fig. 5.
Table 5: The mean of the weighted MMD results over 10 rounds for EDG and EDG (reweighted), each using 5,000 samples. For EDG, all samples follow a uniform distribution. For EDG (reweighted), generated samples of EDG are assigned normalized weights, which are calculated according to Sec. 3.3. The standard deviations are omitted as they are negligibly small.
EDG EDG (reweighted)
Mean value of the reweighted MMD $0.50$ $0.09$
@@
   fc $30\times 256$, ReLU, $2\times$(fc $256\times 256$, ReLU), fc $256\times d$.    fc $(d+31)\times 256$, ReLU,
@@
Final network ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}(\mu,\Sigma)$:
@@
The model architecture of Lennard-Jones
Decoding process Encoding process
Network $a(\zeta;\phi)$: Score network:
   fc $100\times 32$, ReLU, $2\times$(fc $32\times 32$, ReLU), fc $32\times d$.    fc $(d+101)\times 256$, ReLU,
GHD:    fc $256\times 256$, ReLU,
   $Q_{v},T_{v}$:    fc $256\times 100$.
   fc $(2d+1)\times 32$, ReLU, fc $32\times 32$, ReLU, fc $32\times d$.
   $Q_{x},T_{x}$:
   fc $(d+1)\times 32$, ReLU, fc $32\times 32$, ReLU, fc $32\times d$.
Final network $\mathcal{N}(\mu,\Sigma)$:
   fc $d\times 32$, ReLU, fc $32\times 32$, ReLU, fc $32\times d$.
Lennard-Jones The LJ potential is an intermolecular potential which models repulsive and attractive interactions of non-bonding atoms or molecules. The energy is based on the distance of interacting particles $\mathcal{E}^{\mathrm{LJ}}(\cdot)$ and a harmonic potential $\mathcal{E}^{\mathrm{osc}}(\cdot)$ like [klein2023equivariant]
$\mathcal{E}^{\mathrm{LJ}}(x)$ $=$ $\frac{1}{2}\sum_{ij}\left(\left(\frac{1}{d_{ij}}\right)^{6}-\left(\frac{1}{d_{ij}}\right)^{12}\right),$
$\mathcal{E}^{\mathrm{osc}}(x)$ $=$ $\frac{1}{2}\sum_{i}\left\|x_{i}-x_{\mathrm{COM}}\right\|^{2},$
where $d_{ij}=\|x_{i}-x_{j}\|^{2}$ is the Euclidean distance between particles $i$ and $j$ and $x_{\mathrm{COM}}$ refers to the center of mass of the system. The target energy function has the form $\mathcal{E}(\cdot)=\mathcal{E}^{\mathrm{LJ}}(\cdot)+\mathcal{E}^{\mathrm{osc}}(\cdot)$. The system is performed without periodic boundary conditions, which means particles are not wrapped across boundaries, no periodic images or minimum-image convention are used, and interactions are conducted within the finite simulation domain by harmonic confinement. For the experimental results, we evaluate the generated sample with the test sample from [klein2023equivariant]. The test data were generated with MCMC with 1000 parallel chains, where each chain is run for 10,000 steps after a long burn-in phase of 200,000 steps starting from a random generated initial state. The density estimation results of interatomic distances for EDG and BG are plotted in Fig. 3. The results for L2HMC and PIS are presented here in Fig. due to significant deviations from the test data.
Figure 6: Comparison of the test data interatomic distance of LJ13 (left), LJ55 (right) with samples generated from L2HMC and PIS. Results for BG and EDG is ploted in Fig. 3.
Table 6: rESS with mean $\pm$ standard deviation over 10 seeds. For each method, we use 256 samples for estimation.
rESS BG PIS EDG
LJ13 $006\pm 0.002 0.0$$005\pm 0.001$ $\mathbf{0.132\pm 0.048}$
[…]
To evaluate the effectiveness of the sample reweighting strategy described in Sec. 3.3, we employ the relative Effective Sample Size (rESS), defined as
$\mathrm{rESS}=\frac{1}{N}\mathrm{ESS}=\frac{\left(\sum_{n=1}^{N}w(x^{n},z_{0}^{n})\right)^{2}}{N\sum_{n=1}^{N}w(x^{n},z_{0}^{n})^{2}}.$ (17)
Here, $\mathrm{ESS}$ denotes the Effective Sample Size [freeman1966kish], a widely used metric for evaluating the quality of weighted samples produced by importance sampling based on NFs (see, e.g., [wirnsberger2023estimating, schebek2024efficient, jung2024normalizing, coretti2025learning]). The rESS value lies in the interval $(0,1]$ and measures the fraction of effectively independent samples relative to the total sample size $N$. A low rESS indicates that the resulting estimators may suffer from high variance due to poor weight balance.
Ising model The model architecture is shown below, which is similar to that in NeuralRG [ising] without stacking bijectors to form a reversible transformation. It retains the multiscale entanglement renormalization ansatz structure, while we only use one block to update the whole dimensions of the variable directly. To evaluate the efficiency of our proposed architecture, we include a comparative analysis against a standard MLP design in J. The normalizing constant $\log Z$ is approximated according to Sec. 3.3.
@@
   hierarchy network as shown in [ising]    hierarchy network.
@@
Final network ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}(\mu,\Sigma)$:
@@
Appendix I Computation cost
To demonstrate the computational efficiency of our model, we have compared the wall-clock training time and memory usage on the 2D energy (Mog2) task for the following models: the trainable MCMC model L2HMC [mc-5], the VI-based model BG [key-4], and the simulation-based model PIS [diff-6]. For L2HMC and PIS, we adopted the default settings from their respective original papers. The network architectures for BG and EDG are detailed in H. All experiments were conducted on a single NVIDIA RTX 2080 Ti GPU with 11GB of VRAM. The runtime comparison is shown in Fig. 7, and memory usage is reported in Table 7.
Figure 7: Computational time comparison across different methods. Left: the logarithm of the per-epoch training time for each method under a consistent batch size of 128. Right: the evolution of MMD in the Mog2 task with respect to cumulative training time, where all evaluation time has been excluded.
Table 7: Training memory usage of different methods, measured by torch.cuda.memory_allocated.
L2HMC BG PIS EDG
Memory (MB) 17.19 22.41 55.83 16.75
Figure 8: The computational cost of sample reweighting in BG, PIS and EDG. Both PIS and EDG require temporal integration, for which distinct numerical schemes are employed to achieve comparable precision. PIS utilizes torchsde.sdeint with the Stochastic Runge-Kutta (SRK) method and a fixed step size of $T/1000$, while EDG adopts scipy.integrate.solve_ivp using the RK45 method with atol=1e-5, rtol=1e-5.
The results show that EDG reaches competitive accuracy in less training time and with lower memory consumption compared to the other methods. Although the per-epoch training time of BG is shorter than that of EDG, EDG achieves better performance in fewer epochs.
Additionally, we compared the reweighting time cost of three methods that support post-hoc sample reweighting: BG, PIS, and our proposed EDG. The results are presented in Fig. 8. Among them, BG achieves fast computation of the weight function due to the efficient evaluation of the Jacobian determinant in each affine coupling layer. In contrast, PIS and EDG must solve stochastic and ordinary differential equations respectively to compute weights, resulting in higher computational cost. The reweighting time of EDG and PIS is of the same order of magnitude, and EDG is slightly slower. This is mainly because the reweighting in EDG relies on the neural ODE technique, which involves divergence computation that contributes significantly to the time cost. PIS does not require divergence computation during reweighting. It is worth emphasizing that PIS also requires numerical SDE solvers during training, while EDG avoids solving either SDEs or ODEs during training. As shown in Fig. 7, this leads to significantly lower training time for EDG compared to PIS.
Appendix J Ablation study
GHD in Decoder: To elucidate the function of GHD in EDG, we compare the following models on the sampling task of 2D energy functions: a VAE with the Gaussian decoder and encoder, where the means and diagonal covariance matrices are both parameterized by MLPs; a VAE with a GHD-based decoder and MLP-based encoder; EDG without GHD, where the decoder is modeled by MLPs; and the full EDG model. For VAE w/o GHD, the network is composed of 3-layer MLPs, each with 32 units and ReLU activation function. Starting from $z\sim\mathcal{N}(0,I_{d}),\ d=10$, the decoder generates samples $x\sim p_{D}(x;\mu(z;\phi),\Sigma(z;\phi))$. The encoder, as an independent network, follows $p_{E}(z|x)=\mathcal{N}(z;\mu(x;\theta),\Sigma(x;\theta))$. The objective function is the KL divergence of the joint distribution of $z$ and $x$ as